OpenStudy (anonymous):

Consider the function on the interval (0, 2π). x/2 +sin(x) Apply the First Derivative Test to identify the relative extrema. relative maximum (x, y)= relative minimum (x, y) = I believe the derivative is 1/2 -cos(x), but i don't know how to find the max or min. i found the critical numbers: pi/3 & 5pi/3 Please help

OpenStudy (anonymous):

you got the derivative?

OpenStudy (anonymous):

oh yeah,i see that you do have it, but a slight mistake it should be \[\large f'(x)=\frac{1}{2}\color{red}+\cos(x)\]

OpenStudy (anonymous):

if you set that equal to zero and solve, you get \[\cos(x)=-\frac{1}{2}\] so \[x=\frac{2\pi}{3},\frac{4\pi}{3}\]

OpenStudy (anonymous):

the first is a local max, the second is a local min

OpenStudy (anonymous):

@satellite73 that's all? how to find y? would y be 0?

OpenStudy (anonymous):

way beyond my level...